For the reaction


2NH


3

(g)↽−

−⇀

 3H


2

(g)+N


2

(g)

2⁢NH3⁢(g)⁢↽−−⇀ 3⁢H2⁡(g)+N2⁡(g)

the equilibrium concentrations were found to be [NH


3

]=0.250 M

[NH3]=0.250 M, [H


2

]=0.450 M

[H2]=0.450 M, and [N


2

]=0.800 M

[N2]=0.800 M. What is the equilibrium constant for this reaction?


1
Expert's answer
2021-05-08T23:48:56-0400

Ans:-

Reaction is

"2\u2062NH_3\u2062(g)\u2062\\iff 3\u2062H_2\u2061(g)+N_2\u2061(g)"


We have "[NH_3]=0.250 M" "[H_2]=0.450 M" "[N_2]=0.800 M"


"\Rightarrow" Equilibrium constant ("K")"=\dfrac{[H_2]^3\times[N_2]}{[NH_3]^2}"


"\Rightarrow" Equilibrium constant ("K")="\dfrac{(0.450)^3\times(0.800)}{(0.250)^2}=1.1664"


"\therefore" Equilibrium constant ("K")="1.1664"